Company M has issued bonds previously. The bonds have a par value of $1,000 and offer an annual coupon rate of 6%. The bond has 9 years remaining until its maturity date.

Calculate the value (price) of the bond assuming:

The current market interest rate (also called the discount rate, required rate of return, Yield to Maturity) is...

(A) 4%

(B) 8%

(C) 6%
Answer #1

Value of bond = Coupon Amount * PVIFA (YTM, no of years) + Par Value * PVIF (YTM, no of years)

a. The current market interest rate is 4%

Value of bond = (1000 * 6%) * PVIFA (4%, 9 years) + 1000 * PVIF (4%, 9 years)

= 60 * 7.437 + 1000 * 0.703

= 446.22 + 703

= 1149.22

b. The current market interest rate is 8%

Value of bond = (1000 * 6%) * PVIFA (8%, 9 years) + 1000 * PVIF (8%, 9 years)

= 60 * 6.246 + 1000 * 0.5

= 374.76 + 500

= 874.76

c. The current market interest rate is 6%

Value of bond = (1000 * 6%) * PVIFA (6%, 9 years) + 1000 * PVIF (6%, 9 years)

= 60 * 6.8 + 1000 * 0.592

= 408 + 592

= 1000